# Random Latent Exploration for Deep Reinforcement Learning

### v1

@@ 4.1 Illustrative Experiments on FourRoom
4.2 Benchmarking Results on Atari
Having performed illustrative experiments on the FourRoom toy environment, we now evaluate our method on more realistic and challenging tasks. We aim to show that RLE can improve PPO’s overall performance on most tasks.
Setup. We evaluate our method in the well-known Atari benchmark (Bellemare et al., 2013). Following the common practice in Atari (Mnih et al., 2015), the agent perceives a stack of the most recent four $84\times 84$ grayscale frames as inputs and takes discrete actions available in the gamepad of Atari (see Bellemare et al. (2013) for further environment details). For RLE, we chose the feature learned by the value network followed by a randomly initialized linear layer as $\phi$ (used in Equation 2) and set the dimension of the latent vector $\boldsymbol{z}$ as $8$. Note that the randomly initialized linear layer is frozen after initialization. We use the standard PPO hyperparameters provided in (Burda et al., 2019). The hyperparameters and implementation details of all the algorithms and PPO are deferred to Appendix B. For each Atari game (i.e., environment), we train each agent with $5$ different random seeds for $40$ million frames, as prior work (Chen et al., 2022; Bellemare et al., 2016) suggested. However, we trained Montezuma’s Revenge for $200$ million frames since its exploration difficulty is much harder than other Atari games (Burda et al., 2019).
@@ 4.2 Benchmarking Results on Atari
Does RLE improve over the baselines consistently? In addition to the margin of performance improvement on the aggregated score across all games, we demonstrate that RLE results in performance improvement over the baselines with high probability. Following the evaluation protocol suggested in Agarwal et al. (2021), we measure the probability of improvement (POI) between algorithms and their 95% confidence intervals, estimated using the bootstrapping method (DiCiccio & Efron, 1996), and present the results in Figure 5 and Figure 23. Figure 5(a) shows that the lower confidence bound of POI for RLE over each algorithm is above 0.5, indicating that RLE statistically outperforms the other baselines (Agarwal et al., 2021). This means that for a randomly chosen task in Atari, running RLE is likely to yield a higher score than the other baselines, implying that RLE’s performance improvements are consistent and not limited to a few games. Conversely, Figure 23(a) reveals that the POI over PPO for both NoisyNet and RND is below 0.5, suggesting that NoisyNet and RND do not consistently improve over PPO despite having better performance in a few games (see Figure 18). We use a CNN to represent the policy and value function in all Atari experiments, rather than an LSTM as done in (Chen et al., 2022) as the CNN-based architectures were also used in prior work (Burda et al., 2019) and due to its simplicity in implementation. This could explain why PPO outperforms RND over all 57 Atari games in our experiments.
@@
4.3 Benchmarking Results on Isaac Gym
To demonstrate that RLE can improve upon PPO in both discrete and continuous control tasks, we also conducted experiments in IsaacGym (Makoviychuk et al., 2021), a benchmark suite containing numerous continuous control tasks. We compare the performance of RLE to that of PPO and RND as baselines. As RLE and RND use two reward streams with potentially different scales, we use reward normalization for these methods. Thus, we compare against PPO with reward normalization as an additional baseline.
We implemented RLE on top of PPO and trained it with standard PPO hyperparameters in IsaacGym provided in CleanRL, with implementation details provided in Section B.4. We selected both manipulation and locomotion tasks for evaluation and presented the learning curves in Figure 6, with learning curves on all environments presented in Figure 22.
Does RLE improve over baselines in continuous control? The results show that RLE achieves a higher average return than PPO in most tasks, with particularly large performance gains in AllegroHand and ShadowHand, indicating that RLE improves upon PPO in continuous control tasks. In AllegroHand and ShadowHand, the objective is to control a anthropomorphic hand to reorient objects to a target pose. These tasks require more exploration than other continuous control tasks since it takes many steps to achieve the target pose. Additionally, in Cartpole, PPO performance degrades abruptly in the middle of training, while RLE maintains high performance throughout, suggesting that RLE prevents the learning process from collapsing during training.
Does RLE achieve higher overall performance than PPO? To study the overall performance, we measure the probability of improvement of RLE over PPO and RND across 9 different IsaacGym tasks and also present the results in Figure 5 and Figure 23. Figure 5(b) shows that RLE has a statistically significant chance of improving over both PPO and RND in IsaacGym tasks as the lower confidence bound of the POI for RLE over each baseline method is greater than $0.5$. Furthermore, Figure 23(b) shows that out of all considered methods, RLE is the only one with a statistically significant POI over PPO.
We also measure the aggregated normalized return of RLE, PPO, and RND across all $9$ IsaacGym tasks. As the return is positive for PPO in each environment, we normalize runs by dividing by the mean score of PPO in that environment (i.e., normalize PPO to have a score of 1). For further details of this metric, see Section B.4.2. We present the results in Figure 24 and Figure 25, which show that RLE achieves higher aggregate performance than PPO and matches RND in this metric.
@@ 4.4 Ablation Studies
We ran various ablation studies on Atari and IsaacGym environments to explore how different hyperparameters and design choices affect RLE performance.
Latent vector distribution. We investigated the impact of different latent vector distributions on RLE’s performance. Our study involves training RLE with various distributions, including $\mathrm{Uniform}([-0.5,0.5]^{d})$ and isotropic normal $\mathcal{N}(\boldsymbol{0},\mathbf{I}_{d})$ distributions, within a $d$-dimensional space where $d=8$. The detailed implementation is described in Appendix B. The results presented in Figure 7 indicate that RLE performs better than PPO across different latent vector distributions. This suggests that RLE’s efficacy is not significantly affected by the choice of latent vector distribution.
@@
Latent vector dimension. This study explores how robust is RLE to different choices of the dimension $d$ of the latent vector z. We trained RLE for $d\in\{2,8,32,128\}$, where $d=8$ is the dimension used in the results presented in Section 4.2. The outcomes, depicted in Figure 8, demonstrate that RLE is capable of surpassing PPO across all tested dimensions. Although there are slight performance variations between different $d$ values, these differences are subtle, suggesting that RLE’s performance is insensitive to the choice of latent vector dimension $d$.
@@
Feature vectors obtained from random neural network.
In the Atari experiments described in Section 4.2, we used a slow-moving estimate of the CNN features (see Section B.3.1) learned by the value network to compute RLE rewards. This choice of features slightly contributes to improved performance. Figure 20 presents the IQM of the normalized score and the POI of RLE over PPO with and without using the CNN features learned by the value network. The results demonstrate that incorporating value network features leads to a higher POI, while not significantly affecting the IQM. We plot the learning curve in each game for both variants in Figure 21, finding that while performance is broadly similar, there are a few games where the two variants have large differences in performance.
@@
Since RLE relies on neural networks to extract features for computing random rewards $F(s,\boldsymbol{z})$, it’s important to examine how the choice of network architecture affects performance. We investigated the impact of different network architectures for extracting features $\phi(s)$ on the computation of RLE rewards $F(s,\boldsymbol{z})$ in IsaacGym (Makoviychuk et al., 2021), with the IQM of the normalized score in Figure 27 and POI over PPO shown in Figure 26. In our original IsaacGym experiments, we used the value network’s architecture for RLE. In this ablation study, we tested a shallower neural network architecture. The results indicate that RLE with a shallower network still performs well, suggesting that RLE is not highly sensitive to the choice of network architecture.
@@
In Section 3.1, we emphasized the importance of ensuring that RLE’s random reward function $F$ is correlated with states. If the random rewards are not state-dependent, they will act as white noise and will not enhance performance. In this ablation study, we empirically investigate the significance of making random rewards dependent on states. We compare the state-dependent RLE reward $F(s,\boldsymbol{z})$ with white noise rewards sampled from a normal distribution with zero mean and unit variance. This study was conducted in IsaacGym environments, with the results presented in Figure 28. The results demonstrate that white noise rewards significantly degrade performance, indicating that RLE rewards are not merely white noise.
@@ Appendix D Detailed Results and Learning Curves on all Atari Games
  • Aggregated capped human normalized score (described in Section B.3.2) for each of the algorithms (RLE (Ours), PPO, RND and NoisyNet) over all 57 Atari games.
  • An ablation study of how the soft update rule for the feature network affects performance on Atari games taking three metrics into account (scores on individual games, IQM of human normalized score, and probability of improvement over PPO).
@@ Appendix E Detailed Results and Learning Curves on all IsaacGym Tasks
  • The learning curves for PPO and RLE in all 9 IsaacGym tasks that we consider in Figure 22.

### v2

@@ 4.2 Benchmarking Results on Atari
Does RLE improve over the baselines consistently? The aggregated performance improvement across all games reported in the previous Section doesn’t reveal how many games RLE outperforms the baselines. If RLE outperforms baselines on almost all games, where each game can be thought as a different MDP, then we could say that with high-probability RLE will win against other algorithms on new MDPs/tasks. To evaluate this, we measure the probability of improvement (POI; Agarwal et al. (2021)) between algorithms and their 95% confidence intervals, estimated using the bootstrapping method (DiCiccio & Efron, 1996) which is reported in Figure 5 and Figure 23 (in Appendix). Figure 5(a) shows that the lower confidence bound of POI for RLE over each algorithm is above 0.5, indicating that RLE statistically outperforms the other baselines (Agarwal et al., 2021). This means that for a randomly chosen task in Atari, running RLE is likely to yield a higher score than the other baselines, implying that RLE’s performance improvements are consistent and not limited to a few games.
Conversely, Figure 23(a) reveals that the POI over PPO for both NoisyNet and RND is below 0.5, suggesting that NoisyNet and RND do not consistently improve over PPO despite having better performance in a few games (see Figure 18). We use a CNN to represent the policy and value function in all Atari experiments, rather than an LSTM as done in (Chen et al., 2022) as the CNN-based architectures were also used in prior work (Burda et al., 2019) and due to its simplicity in implementation. This could explain why PPO outperforms RND on average over the 57 Atari games in our experiments.
@@
4.3 Evaluation on Isaac Gym
To demonstrate that RLE can improve upon PPO in both discrete and continuous control tasks, we also conducted experiments in IsaacGym (Makoviychuk et al., 2021), a benchmark suite containing numerous continuous control tasks. We compare the performance of RLE against PPO and RND. As RLE and RND both use a secondary reward function with potentially different scale from the task reward function, we use reward normalization for these methods (Burda et al., 2019). Thus, we compare against PPO with reward normalization as an additional baseline.
We implemented RLE on top of PPO and trained it with standard PPO hyperparameters in IsaacGym provided in CleanRL, with implementation details provided in Section B.4. The performance of four representative tasks and all the tasks is reported in Figure 6 and Figure 22 (in Appendix) respectively.
Does RLE improve over baselines in continuous control? The results show that RLE achieves a higher average return than PPO in most tasks, with particularly large performance gains in AllegroHand and ShadowHand, indicating that RLE improves upon PPO in continuous control tasks. In AllegroHand and ShadowHand, the objective is to control an anthropomorphic hand to reorient objects to a target pose. These tasks require more exploration than other continuous control tasks since it takes many steps to achieve the target pose. To study the overall performance, we measure the probability of improvement of RLE over PPO and RND across nine different IsaacGym tasks and also present the results in Figure 5 and Figure 23. Figure 5(b) shows that RLE has a statistically significant chance of improving over both PPO and RND in IsaacGym tasks as the lower confidence bound of the POI for RLE over each baseline method is greater than $0.5$. Furthermore, Figure 23(b) shows that out of all considered methods, RLE is the only one with a statistically significant POI over PPO. To estimate the average performance difference across environments, we also measure the aggregated normalized return of RLE, PPO, and RND across all IsaacGym tasks. As the return is positive for PPO in each environment, we normalize runs by dividing by the mean score of PPO in that environment (i.e., normalize PPO to have a score of 1). For further details of this metric, see Section B.4.2. We present the results in Figure 24 and Figure 25, which show that RLE achieves higher aggregate performance than PPO and matches RND in this metric.
@@ 4.4 Ablation Studies
We ablated various design choices in RLE using both the Atari and IsaacGym benchmarks.
Effect of latent vector distribution. We investigated the impact of different latent vector distributions on RLE’s performance by choosing different sampling strategies including $\mathrm{Uniform}([-0.5,0.5]^{d})$ and isotropic normal $\mathcal{N}(\boldsymbol{0},\mathbf{I}_{d})$ distributions, within a $d$-dimensional space where $d=8$. The detailed implementation is described in Appendix B. The results presented in Figure 7 indicate that RLE performs better than PPO across different latent vector distributions. This suggests that RLE’s efficacy is not significantly affected by the choice of latent vector distribution.
@@
Effect of latent vector dimension. This study explores how robust is RLE to different choices of the dimension $d$ of the latent vector z. We trained RLE for $d\in\{2,8,32,128\}$, where $d=8$ is the dimension used in the results presented in Section 4.2. The outcomes, depicted in Figure 8, demonstrate that RLE can surpass PPO across all tested dimensions. Although slight performance variations exist between different $d$ values, these differences are subtle, suggesting that RLE’s performance is relatively insensitive to the choice of latent vector dimension $d$.
@@
Since RLE relies on neural networks to extract features for computing random rewards $F(s,\boldsymbol{z})$, it is important to examine how the choice of network architecture affects performance. We investigated the impact of different network architectures for extracting features $\phi(s)$ on the computation of RLE rewards $F(s,\boldsymbol{z})$ in IsaacGym (Makoviychuk et al., 2021). We test only IsaacGym as Atari requires a different architecture to handle the image observation space, leaving no clear comparison with the architecture variations investigated here. We display the IQM of the normalized score in Figure 27 and POI over PPO in Figure 26. In our original IsaacGym experiments, we used the value network’s architecture for RLE. In this ablation study, we tested a shallower neural network architecture. The results indicate that RLE with a shallower network still performs well, suggesting that RLE is not highly sensitive to the choice of network architecture.
@@
In Section 3.1, we emphasized the importance of ensuring that RLE’s random reward function $F$ is state-dependent. We compare the state-dependent RLE reward $F(s,\boldsymbol{z})$ with white noise rewards sampled from a normal distribution with zero mean and unit variance. We conduct this study in IsaacGym environments, with the results presented in Figure 28. The results demonstrate that white noise rewards significantly degrade performance, indicating that RLE rewards are not merely white noise.
@@ Appendix D Detailed Results and Learning Curves on all Atari Games
  • Aggregated capped human normalized score (described in Section B.3.2) for each of the algorithms (RLE (Ours), PPO, RND and NoisyNet) over all 57 Atari games in Figure 19
  • An ablation study of how the soft update rule for the feature network affects performance on Atari games taking three metrics into account (scores on individual games, IQM of human normalized score, and probability of improvement over PPO). In the Atari experiments described in Section 4.2, we used a slow-moving estimate of the CNN features (see Section B.3.1) learned by the value network to compute RLE rewards. This choice of features slightly contributes to improved performance. Figure 20 presents the IQM of the normalized score and the POI of RLE over PPO using the CNN features learned by the value network vs. using the randomly initialized CNN features. The results demonstrate that incorporating value network features leads to a higher POI, while not significantly affecting the IQM. We plot the learning curve in each game for both variants in Figure 21, finding that while performance is broadly similar, there are a few games where the two variants have significant differences in performance.
@@ Appendix E Detailed Results and Learning Curves on all IsaacGym Tasks
  • The learning curves for PPO and RLE in all 9 IsaacGym tasks that we consider in Figure 22. Note that in Cartpole, PPO performance degrades abruptly in the middle of training, while RLE maintains high performance throughout, suggesting that RLE prevents the learning process from collapsing during training.